OpenStudy (anonymous):

Find the Laplace Transform of the given function; a and b are real numbers and n is a positive integer.

OpenStudy (anonymous):

\[\huge f(t)=t e^{at} \sin(bt)\] Here's what I have.. \[\large \text{Let}~~g(t)=e^{at} \sin(bt)\]\[\large G(s)=\mathcal{L}[g(t)]=\frac{b}{(s-a)^2+b^2}\]\[\text{Then}~~F(s)=G'(s)\](which I haven't done yet.. just wanted to check if I'm on the right track here..)

OpenStudy (anonymous):

@ganeshie8 X)

ganeshie8 (ganeshie8):

are you trying to use convolution ? @dan815 is online

ganeshie8 (ganeshie8):

convolution may not be needed here, i think your method works nicely

OpenStudy (anonymous):

I'm not sure.. I found the solution to a similar problem for \(t^2 \sin(bt)\) and they followed the same process using this property: \[\mathcal{L}[t^n f(t)]=(-1)^n \cdot \frac{d}{ds^n}F(s)\]

OpenStudy (anonymous):

which.. i forgot the -1

ganeshie8 (ganeshie8):

\[\mathcal{L}\{t e^{at} \sin(bt)\} = -\partial_s\mathcal{L}\{e^{at}\sin (bt)\}\]

OpenStudy (anonymous):

Is that partial derivative with respect to s?

ganeshie8 (ganeshie8):

next you want to use \[\mathcal{L}\{e^{at}f(t)\} = F(s-a)\] right ?

ganeshie8 (ganeshie8):

yes it is w.r.t s

OpenStudy (anonymous):

Just checkin! and yeah

ganeshie8 (ganeshie8):

Looks good to me!

ganeshie8 (ganeshie8):

@dan815 may show you another way to work this using convolution

OpenStudy (anonymous):

@ganeshie8 Okay, cool! And just for future reference.. will there be a time when this property can't be used for problems that may *look* similar?

OpenStudy (anonymous):

Also, my final answer is: \[\huge F(s)=\frac{ 2b(s-a) }{ ((s-a)^2+b^2)^2 }\]

ganeshie8 (ganeshie8):

hey sorry there is some issue with notifications, below property works always \[\mathcal{L}\{e^{at}f(t)\} = F(s-a)\] ofcourse laplace transform of f(t) need to exist for above property to make any sense

OpenStudy (anonymous):

@ganeshie8 It's all good and sweet! X) Thanks